# A Poisson-Guided Decomposition Network for Extreme Low-Light Image Enhancement

### v1

@@ Abstract
Low-light image denoising and enhancement are challenging, especially when traditional noise assumptions, such as Gaussian noise, do not hold in majority. In many real-world scenarios, such as low-light imaging, noise is signal-dependent and is better represented as Poisson noise. In this work, we address the problem of denoising images degraded by Poisson noise under extreme low-light conditions. We introduce a light-weight deep learning-based method that integrates Retinex based decomposition with Poisson denoising into a unified encoder-decoder network. The model simultaneously enhances illumination and suppresses noise by incorporating a Poisson denoising loss to address signal-dependent noise. Without prior requirement for reflectance and illumination, the network learns an effective decomposition process while ensuring consistent reflectance and smooth illumination without causing any form of color distortion. The experimental results demonstrate the effectiveness and practicality of the proposed low-light illumination enhancement method. Our method significantly improves visibility and brightness in low-light conditions, while preserving image structure and color constancy under ambient illumination.
@@ I Introduction
Images captured under low-light conditions often suffer from high noise levels, poor contrast, and distorted illumination, resulting in significant degradation in visual quality and make it difficult to extract meaningful information from images. This issue is prominently affecting real-time systems, such as autonomous driving [1], object detection [2] where accurate and clear image representation is essential for decision making. Therefore, efficient computational techniques are particularly important for reducing the loss of system performance caused by low-light images.
@@
Conventional image enhancement techniques such as [24, 29, 12, 30], [13, 31], often assume that noise is independent of the signal and is modeled as zero-mean additive Gaussian noise [32]. However, in practical scenarios such as low-light imaging the noise is typically signal-dependent and is better represented by Poisson statistics [33, 34]. The noise in extreme low-light images is dominated by Poisson statistics, which differs from the Gaussian noise assumption made by many standard denoising algorithms [35, 36]. This distinction makes conventional methods less effective in handling the noise typically found in low-light environments [37]. The illumination in such images is often distorted due to extreme low illumination, making it challenging to recover true scene details. As a result, the performance of most existing techniques is limited under low-light conditions, where the signal-to-noise ratio (SNR) is low and the photon counts are limited [38, 39, 40].
@@
In this paper, we propose a deep learning method for illumination recovery and enhancement of images corrupted by Poisson noise due to poor lighting conditions. In particular, we introduce a unified encoder-decoder network which integrates Retinex based decomposition with Poisson denoising. Our encoder-decoder network learns an effective decomposition process while ensuring consistent reflectance and smooth illumination components. Our proposed modified Retinex decomposition has an additional component representing Poisson noise at each pixel. Our novelty is that we have a dedicated loss function in our deep learning architecture to suppress the noisy component in the output/enhanced image.
@@
The rest of the paper is organized as follows. In Section II, we introduce the Poisson denoising problem and our proposed Poisson guided decomposition of extreme low-light image. We present our proposed deep learning network in Section III and experimental results in Section IV. Finally, we conclude with a summary in Section V.
@@
II Poisson denoising problem
The primary source of noise in digital images arises during image acquisition or transmission. In the case of low-light images with poor exposure, noise is predominantly due to a low signal-to-noise ratio (SNR) and low photon count: fewer photons reaching the sensor limit the number of photons collected by a pixel. This results in increased uncertainty, as described by the Poisson distribution, which models the occurrence of independent random events such as photon arrivals. Consequently, Poisson noise, also known as shot noise, becomes the dominant form of noise in such conditions.
Let $X$ denote a noisy image produced by a sensor. The objective of denoising is to recover a clean latent image $Y$. In extreme low-light imaging, the noise is dominated by Poisson noise. Given the true value $Y(i)$ of the $i$-th pixel expressed in the number of photons, the corresponding value of the observed pixel $X(i)$ is an independent Poisson distributed random variable;
$P(X(i)=n)=\frac{\lambda^{n}e^{-\lambda}}{n!},$ (1)
where $\lambda>0$ is both mean and variance [41]. In defining the Poisson statistics for each pixel of the low-light image, it is assumed that the number of arrivals during non-overlapping time intervals are independent random variables and there exists a positive quantity $\lambda$ which is the contribution of a photon to the intensity of the pixel in $\Delta t$ duration and satisfying the postulates for the distribution of Poisson for each value of the pixels [41].
II-A Our Poisson-Guided Image Decomposition
@@
It is well known that low-light digital images are typically captured and stored in RGB format [42], and since Poisson noise in the sensor affects all three color channels. This can lead to color distortion [43], [44] during image enhancement. To preserve color constancy under ambient illumination, we consider to decompose the captured color image $X$ as following:
$X^{c}(i)=L(i)\circ R^{c}(i)\circ N^{c}(i),\quad c\in\{r,g,b\},$ (2)
@@
where $\circ$ indicates element-wise multiplication at each pixel $i$; $L$, $R$, and $N$ are illumination, reflectance, and noisy components respectively. Poisson noise is neither additive nor multiplicative, but it is signal-dependent. However, it can be formulated as a multiplicative component of the noise in the pixel values. To preserve color constancy and prevent color shifts during enhancement, we formulate a Poisson-guided image decomposition framework in which an image is decomposed into element-wise multiplication of three components: illumination $L$, reflectance ($R$) and noisy ($N$).
Figure 2: Visual comparison with other state-of-the-art low-light enhancement methods: (a) low-light input image, (b) LIME [13], (c) deep Retinex [24], (d) SRIE [45], (e) zero-DCE [46], (f) Proposed, (g) Ground Truth. Our method produces superior quality output image without any color distortion.
III Proposed Method
III-A Network Architecture
@@
We propose a light-weight convolutional encoder-decoder network shown in Figure 1.
III-A1 Initial Convolution Layer
@@
The first layer is a convolutional layer with 3 input channels (RGB input images patches) of size $3\times 128\times 128$ and $64$ output channels, utilizing a $3\times 3$ kernel size. This layer extracts basic low-level features from the input image, such as edges and textures.
III-A2 Encoder-Decoder Layers
@@
The encoder consists of three consecutive convolutional layers with strides of two. Each convolutional layer is followed by a ReLU activation function to introduce nonlinearity for better feature representation. The decoder consists of three deconvolutional layers, each taking input from both the corresponding encoder layer (via concatenation) and the previous decoder layer. This enables the network to recover spatial resolution while combining high-level and low-level features for better image reconstruction.
III-A3 Image Reconstruction (Output Layers)
The network is divided into three parallel output layers to predict three distinct components: illumination map (via sigmoid activation function), reflectance map (via sigmoid activation) and noise (using a tanh activation function). Each output is resized to match the original input dimensions. Finally, the output enhanced image is the product of illumination map and reflectance map.
III-B Loss Function
@@
The loss function is designed to optimize the performance of the model by incorporating multiple components that guide the network to perform image reconstruction, noise suppression, and smoothness regularization. Our proposed loss function is as follows:
@@ III-B1 Reconstruction Loss
$\mathcal{L}_{rec}=\parallel X-(R\circ L\circ N)\parallel_{1}$ (4)
which measures the fidelity between the original and reconstructed images while providing robustness to outliers and encouraging sparsity in the reconstruction error.
III-B2 Retinex Estimation Loss
@@
where $\gamma$ is a hyperparameter. Further, loss due to illumination-reflectance decomposition error is given by:
$\mathcal{L}_{decom}=\parallel Y-R\circ L\parallel_{1}.$ (6)
@@
whereas, $\beta$ represents the coefficients for balanced structure and smoothness of illumination map, which is defined as $10$ in our experiments and $\Delta R$ represents gradient of reflectance in both x and y direction, $\Delta L$ represents gradient of Illumination in both x and y direction. The term exponential is used to protect against structural loss in the illumination map, and the gradient ensures the smoothness of the illumination.
III-B3 Noise Loss
The loss term works as a constraint to extract the noisy component from the low-light image:

### v2

# PIRL: Poisson Informed Retinex Network for Extreme Low-Light Image Enhancement

@@ Abstract
Low-light image denoising and enhancement is a critical task in computer vision, with applications in photography, surveillance, autonomous driving, and medical imaging. In such conditions, images often suffer from severe degradation due to low photon counts, making both enhancement and denoising particularly challenging. In many real-world scenes of low-light imaging, noise, which is in majority signal-dependent, is better represented as Poisson noise. In this work, we address the problem of denoising images degraded by Poisson noise under extreme low-light conditions. We propose a novel Poisson noise modeling for extreme low-light images, which is then directly integrated into a trainable deep learning network for low-light image enhancement. We propose an end-to-end deep learning method based on Poisson noise informed retinex decomposition. Our proposed method simultaneously enhances illumination and suppresses noise by incorporating a Poisson denoising loss that effectively addresses signal-dependent noise. Without any prior requirement for reflectance and illumination, the network learns an effective decomposition process while ensuring consistent reflectance and smooth illumination without introducing any form of color distortion. The experimental results demonstrate the effectiveness and practicality of the proposed low-light illumination enhancement method. Our method significantly improves visibility and brightness in low-light conditions, while preserving image structure and color constancy under ambient illumination.
@@ I Introduction
Images captured under low-light conditions often suffer from high noise levels, poor contrast, and distorted illumination, which result in significant degradation in the visual quality and make it difficult to extract meaningful information from the images. This issue is prominently affecting real-time systems, such as surveillance, autonomous driving [1], object detection [2] where accurate and clear image representation is essential for decision making. Therefore, efficient computational techniques are particularly important for reducing the loss of system performance caused by low-light images.
@@
Conventional image enhancement techniques based on Retinex decomposition such as [24], [13, 29], often assume that noise is independent of the signal and is modeled as zero-mean additive Gaussian noise [30]. Hence all conventional methods undergoes enhancement and then denoising of images, handling both the tasks as individual and independent of each other. However, in practical scenarios such as low-light imaging the noise is typically signal-dependent and hence is better represented by Poisson statistics [31, 32]. The noise in extreme low-light images is dominated by Poisson statistics, which differs from the Gaussian noise assumption made by many standard denoising algorithms [33, 34]. This distinction makes conventional methods less effective in handling the noise typically found in low-light environments [35]. The illumination in such images is often distorted due to extreme low illumination, making it challenging to recover true scene details. As a result, the performance of most existing techniques is limited under low-light conditions, where the signal-to-noise ratio (SNR) is low and photon counts are limited [36, 37, 38].
@@
In this paper, we propose a Poisson-aware loss function for illumination recovery and enhancement of images corrupted by Poisson noise due to poor lighting conditions. In particular, we introduce a unified light-weight encoder-decoder network which integrates Retinex based decomposition with Poisson denoising. Our encoder-decoder network learns an effective decomposition process while ensuring consistent reflectance and smooth illumination components. Our proposed modified Retinex decomposition has an additional component representing Poisson noise at each pixel. Our novelty is that we have a dedicated loss function in our deep learning architecture to suppress the noise component in the output/enhanced image.
  1. 1.
    We propose a Poisson noise modeling, which is then directly integrated into a trainable deep learning network for low-light image enhancement. By strengthening Retinex decomposition with a physics driven solution, we effectively address the challenges of insufficient information and severe feature degradation in single-image low-light conditions.
  2. 2.
    By incorporating a dedicated noise estimation branch and formulating a noise-aware loss, we enable joint learning of enhancement and denoising in an end-to-end manner. Our composite loss function jointly performs enhancement and denoising of low-light images.
  3. 3.
    We extensively experimented on the recovery of images under extreme low-light conditions of varied severity of lack of illumination. Handling the complexities of signal-dependent noise while preserving structure and color consistency, we are able to achieve superior performance under extreme low-light conditions.
@@
The rest of the paper is organized as follows. In Section III, we introduce the Poisson denoising problem and our proposed Poisson guided decomposition of extreme low-light image. We present our proposed deep learning network in Section IV and the experimental results in Section V. Finally, we conclude with a summary in Section VI.
II Related Works
II-A Classical Methods
In low-light conditions, image quality is often constrained by the optical signal attenuation, resulting in a significant degradation of the signal-to-noise ratio and a notable increase in noise. Early enhancement methods such as histogram equalization and gamma correction aimed to stretch the dynamic range of low-light images. These methods are fast but often cause artifacts like over-enhancement and loss of detail. Adaptive versions like brightness preserving dynamic histogram equalization [3, 4, 5] tried to address these issues but lacked robustness in extreme darkness. Inspired by the human visual system, Retinex theory [8] models an image as the product of reflectance and illumination. Early works like SSR and MSR [10], [9] used handcrafted filters, while later methods integrated bilateral filtering [11] or bright-pass filtering [12]. However, these models typically assume ideal lighting and do not address noise explicitly. Based on Retinex Theory, LIME [13] enhances low-light images by estimating a structure-aware illumination map and applying pixel-wise amplification based on the estimated illumination. While this strategy improves visibility, the method does not incorporate an explicit denoising component within its core algorithm. As a result, it often amplifies the noise inherently present in low-light images. Although the authors recommend using an external denoiser such as BM3D [25] as a post-processing step, this decoupled approach lacks adaptability to varying noise levels and image structures.
II-B Deep Learning Methods
Deep learning-based methods [39, 40, 41, 42, 43] have significantly advanced low-light image enhancement by learning to restore visibility and contrast directly from data. Deep models such as RetinexNet [24] use learned decomposition followed by denoising filters. However, the denoising step is handled externally using BM3D [25], which assumes additive Gaussian noise and is decoupled from the learning process, making it suboptimal for signal-dependent noise encountered in real-world low-light scenarios. Zero-DCE [44] proposed a zero-reference framework that learns pixel-wise brightness adjustment curves without paired supervision. While efficient, it lacks explicit modeling of structural information or noise, often resulting in overexposed or color-distorted outputs under extremely poor illumination. Similarly, EnlightenGAN [45] uses adversarial learning to produce visually appealing results, but it tends to hallucinate textures and is prone to generating inconsistent enhancements.
Recent approaches such as Transformer-GAN [27] and RawFormer [28] employ vision transformers to capture long-range dependencies, improving structure preservation and global enhancement. Despite their strong performance, these models are computationally expensive and not well-suited for real-time or low-power applications. On the other hand, lightweight models [26, 46] aim to reduce complexity using shallow networks or knowledge distillation. However, they often sacrifice enhancement quality and provide limited noise suppression, particularly under extreme low-light where signal-to-noise ratios are very low. While many existing low-light enhancement methods operate under the assumption of additive white Gaussian noise (AWGN), real-world low-light images are primarily degraded by Poisson noise, which is signal-dependent in nature. This is particularly prominent in scenarios with low photon counts, where each pixel’s intensity corresponds to a discrete number of photon arrivals.
II-C Methods for Extreme Low-light Images
To better address this, several recent methods have begun incorporating Poisson noise models into their frameworks. For instance, Kong et al.[47] proposed a Poisson-aware Retinex model that integrates noise-aware constraints into the decomposition of an image into illumination and reflectance. Their method formulates a joint optimization problem that regularizes illumination estimation while accounting for signal-dependent noise. While this approach improves performance under moderate low-light settings, it relies on handcrafted priors and solves the problem through iterative optimization. This makes it computationally inefficient and less suitable for real-time or end-to-end training.
Similarly, Huang et al. [48] introduced a quaternion-screened Poisson equation framework for low-light image enhancement. They extended the screened Poisson equation to quaternion space to jointly process color channels, allowing more consistent illumination correction and smoother gradient preservation. Although their method demonstrates improved color fidelity and brightness restoration, it does not perform explicit denoising and is primarily focused on the enhancement of structural components. Additionally, like other model-based approaches, it lacks learning adaptability and does not support data-driven optimization or joint image reconstruction.
III Poisson denoising problem and Beyond
The primary source of noise in digital images arises during image acquisition or transmission. In the case of low-light images with poor exposure, noise is predominantly due to a low signal-to-noise ratio (SNR) and low photon count—fewer photons reaching the sensor limit the number of photons collected by a pixel. This results in increased uncertainty, as described by the Poisson distribution, which models the occurrence of independent random events such as photon arrivals. Consequently, Poisson noise, also known as shot noise, becomes the dominant form of noise in such conditions.
Let $Y$ denotes a noisy image produced by a sensor. The objective of denoising is to recover a ground-truth clean image $X$ observed by the sensor. In extreme low-light imaging, the noise is dominated by Poisson noise. Given the true value $X(i)$ of the $i$-th pixel expressed in the number of photons, the corresponding value of the observed pixel $Y(i)$ is an independent Poisson distributed random variable;
$P(Y(i)=n|X(i)=\lambda)=\frac{\lambda^{n}e^{-\lambda}}{n!},$ (1)
where $\lambda>0$ is both mean and variance [49]. In defining the Poisson statistics for each pixel of the low light image, it is assumed that the number of arrivals during non-overlapping time intervals are independent random variables. There exists a positive quantity which is the contribution of a photon to the intensity of the pixel in $\Delta t$ duration and satisfying the postulates for the distribution of Poisson for each value of the pixels [49].
@@ III Poisson denoising problem and Beyond
III-A Physics of Poisson Noise via Multiplicative Decomposition
@@
It is well known that low-light digital images captured by cameras suffer from high levels of Poisson noise. These images are typically captured and stored in RGB format [50], and since Poisson noise is signal-dependent, the random nature of photon arrivals at the sensor affects all three color channels. The color image we see is a linear combination of the RGB channels obtained via the Bayer pattern, which can lead to color shifts and failure to preserve color constancy [51], [52] during image enhancement if denoising is not properly addressed [53]. Performing denoising after enhancement is also ineffective in such cases, as it can further degrade the image. Therefore, to preserve color constancy under ambient illumination, it is crucial to apply denoising alongside enhancement.
Existing methods for Poisson denoising often employ a Bayesian framework, treating it as an ill-posed problem to estimate both the illumination and reflectance components from low-light images. Typically, these methods adopt the maximum a posteriori probability (MAP) approach to estimate a noise-free image [34], [54]. Poisson noise is neither purely additive nor multiplicative, but it is signal-dependent. In extreme low-light situations, Poisson noise can become so dominant that the entire observed pixel can be considered Poisson-corrupted, and it can be formulated as a multiplicative component of the noise in the pixel values.
To address this and preserve both color constancy and prevent color shifts, we have formulated a Poisson-aware Retinex model for processing extreme low-light images. As it is independent random process, so individual Poisson corrupted pixel is represented as
$Y(i)=Poisson(I_{i}),\forall i,$ (2)
where $i$ is the index of pixels in the image. From (1) we get
$Y(i)=\frac{(X(i))^{n}e^{-X(i)}}{n!},\forall i.$ (3)
Then, we expand as follows:
$Y(i)=X(i)\frac{(X(i))^{(n-1)}e^{-X(i)}}{n!},\forall i.$ (4)
Thus, the corrupted low-light image is now expressed as product of the noise-fee image $X$ and a component of Poisson noise expressed as:
$N(i)=\frac{(X(i))^{n-1}e^{-X(i)}}{n!},\forall i.$ (5)
Finally, the observed image at the $i$-th pixel is given by
$Y(i)=X(i)N(i),\forall i.$ (6)
So, the low-exposure image $Y$ is represented as element-wise multiplication of the sufficiently exposed image $X$ and the multiplicative component of Poisson noise $N$.
Recalling the retinex theory, an image can be decomposed into element-wise multiplication of illuminance $L$ and reflectance $R$. Therefore,
$X=L\circ R$ (7)
and
$Y=L\circ R\circ N.$ (8)
Finally, by using our hypothesis of Poisson-aware retinex method, a low light image can be decomposed into element wise multiplication of three componets that are multiplicative components of noise $N$, reflectance $R$ and illuminance $L$. Here $\circ$ refers element-wise multiplication of two matrices.
III-B Poisson Noise Guided Color Image Decomposition
It is well known that low-light digital images are typically captured and stored in RGB format [50], and since Poisson noise at the sensor affects all three color channels. This can lead to color shifts and failure to preserve color constancy [51], [52] during image enhancement. To preserve color constancy under ambient illumination, we consider to decompose the captured color image $X$ as following:
$Y^{c}(i)=L(i)\circ R^{c}(i)\circ N^{c}(i),\quad c\in\{r,g,b\},$ (9)
@@ III-B Poisson Noise Guided Color Image Decomposition
where $\circ$ indicates element-wise multiplication at each pixel $i$; $L$, $R$, and $N$ are illumination, reflectance, and noise components respectively. Poisson noise is neither additive nor multiplicative, but it is signal-dependent. However, it can be formulated as a multiplicative component of the noise in the pixel values. To preserve color constancy and prevent color shifts during enhancement, we formulate a Poisson guided image decomposition framework where an image is decomposed into the element-wise multiplication of three components: illumination $L$, reflectance ($R$), and noise ($N$).
IV Proposed Method
IV-A Network Architecture
@@
We propose a convolutional encoder-decoder architecture that is shown in Figure 1. The key components are as follows:
IV-A1 Initial Convolution Layer
@@
The first layer is a convolutional layer with 3 input channels (RGB input image patches) of size $3\times 128\times 128$ and $64$ output channels, utilizing a $3\times 3$ kernel size. This layer extracts basic low-level features from the input image, such as edges and textures.
IV-A2 Encoder-Decoder Layers
@@
The encoder consists of three consecutive convolutional layers with strides of two. We extract hierarchical features at multiple scales by progressively downscaling the features in each cascaded convolutional layer. Each such layer is followed by a ReLU activation function to introduce nonlinearity for better feature representation.
In the decoder stage, we perform feature concatenation via skip connections. This allows for the propagation of fine-grained features from the encoder to the decoder. The feature concatenation operation results in better preservation of details during reconstruction. The decoder consists of three deconvolutional layers. Each of these layers takes input from both the corresponding encoder layer (via concatenation) and the previous decoder layer. This enables the network to recover spatial resolution while combining high-level and low-level features for better image reconstruction.
Figure 2: Schematic of the loss function in our proposed method Poisson Informed Retinex of Low-light images (PIRL).
(a) Input (b) LIME (c) Deep-Retinex (d) Zero-DCE (e) Proposed
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $1$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $2$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $3$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $4$.
Figure 3: Visual comparison of low-light enhancement results for Scene 1: AC image. We compare with recent low-light enhancement methods: (a) low-light input image, (b) LIME, (c) deep-retinex, and (e) Proposed. Notice that our method achieves better quality in the enhanced image than the other methods shown here.
(a) LIME.
(b) Deep-retinex.
(c) Zero-DCE.
(d) PIRL (proposed).
Figure 4: Plot of histograms of RGB channels of recovered AC image at low-light level 3 in Figure 3. An image with good color consistency exhibits very similar histograms for RGB colors. Notice that our method gives the best color consistency resulting no color distortion unlike other methods.
(a) Input (b) LIME (c) Deep-Retinex (d) Zero-DCE (e) Proposed
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $1$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $2$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $3$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Low-light level $4$.
Figure 5: Visual comparison of low-light enhancement results on the Door image:: (a) low-light input image, (b) LIME, (c) deep-retinex, (d) zero-DCE, and (e) Proposed method PIRL. Our method achieves significantly better visual quality in the enhanced image than the other methods shown.
IV-A3 Image Reconstruction (Output Layers)
Our network produces three outputs through three separate branches. The pre-final output of the first branch (top output image shown in Figure 1) is passed through a sigmoid activation to constrain the values in the range $[0,1]$. This output eventually generates a single-channel image that represents the illumination map. The second output is a three-channel image representing the reflectance map, also followed by a sigmoid activation. The third output represents the noise map consisting of three channels. In this case, we apply a tanh activation function to allow negative and positive values to represent noise characteristics.
Notice that we estimate the Poisson noise (three-channel) separately under the assumption of extreme low-light imaging. It is important to note the novel retinex decomposition modeling in this work.
IV-B Loss Function
@@
We propose a new loss function to efficiently train the proposed deep learning model. We optimize the training of our model by incorporating multiple components that guide the network performing image reconstruction, noise suppression, and smoothness regularization. Our loss function consists of multiple losses as follows:
@@ IV-B1 Reconstruction Loss
$\mathcal{L}_{rec}=\parallel Y-(R\circ L\circ N)\parallel_{1}.$ (11)
Reconstruction loss measures the fidelity between the original and reconstructed images. This provides robustness to outliers and encourages sparsity in the reconstruction error.
IV-B2 Retinex Estimation Loss
@@
where $\gamma$ is a hyperparameter that needed to be adjusted based on learning. In addition,
$\mathcal{L}_{decom}=\parallel X-R\circ L\parallel_{1}.$ (13)
Unlike the dark channel prior in the classical dehazing algorithm, or from the bright channel prior incorporated to impose constraints on the illumination map, our proposed deep network decomposes an image into $(R,L,N)$ without any prior for $R$ and $L$. However, we attempt to mitigate the noise component in the enhanced image by applying the loss $\mathcal{L}_{decom}$ in (13).
@@
whereas, $\beta$ represents the coefficients for balanced structure and smoothness of illumination map, which is defined as $10$ in our experiments and $\Delta R$ represents gradient of reflectance in both x and y direction, $\Delta L$ represents gradient of Illumination in both x and y direction. The term exponential is employed to safeguard against structural loss in the illumination map, and gradient ensures the smoothness of the illumination. An ideal illumination map should retain structural details from the input image while maintaining smoothness in texture. To achieve this, we introduce a structure-preserving smoothness loss $L_{ret2}$, inspired by the Total Variation (TV) [55], which is used to weight the gradient of the reflectance map.
IV-B3 Noise Loss
The loss works as a constraint to extract the noisy component and mitigate noise in from the low light image;